Problem 16-12 Calculating WACC [LO1] Blitz Industries has a debt-equity ratio of 1.6. Its WACC is 7.8 percent, and its cost of debt is 5.5 percent. The corporate tax rate is 21 percent. 

a. What is the company’s cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

b. What is the company’s unlevered cost of equity capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

c-1. What would the cost of equity be if the debt-equity ratio were 2? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

c-2. What would the cost of equity be if the debt-equity ratio were 1.0? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

c-3. What would the cost of equity be if the debt-equity ratio were zero? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
D/A = D/(E+D)
D/A = 1.6/(1+1.6)
=0.6154
E/A = 1-D/A
=1-0.6154
=0.3846
a
WACC = Levered cost of equity*E/A+Cost of debt*(1-tax rate)*D/A
0.078= Levered cost of equity*0.3846+0.055*(1-0.21)*0.6154
Levered cost of equity =-0.78%
b
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
-0.00781713947990544 = Unlevered cost of equity+1.6*(Unlevered cost of equity-0.055)*(1-0.21)
Unlevered cost of equity = 2.73
C-1
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 2.73+2*(2.73-0.055)*(1-0.21)
Levered cost of equity = 6.96
C-2
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 2.73+1*(2.73-0.055)*(1-0.21)
Levered cost of equity = 4.84
C-3
Levered cost of equity = Unlevered cost of equity+D/E*( Unlevered cost of equity-cost of debt)*(1-tax rate)
Levered cost of equity = 2.73+0*(2.73-0.055)*(1-0.21)
Levered cost of equity = 2.73
Answer #2

the correct answer should be: 

a.Cost   of equity13.33% 
b.Unlevered   cost of equity8.96% 
c-1.Cost   of equity14.42% 
c-2.Cost   of equity11.69%
c-3.Cost of equity8.96%